If the silver phosphate solubility product –
Ag3PO4 for pKs = 18.2; the concentration of silver in a
saturated solution will be:





1
Expert's answer
2022-03-07T03:33:04-0500

Solution:

silver phosphate = Ag3PO4


pKs(Ag3PO4) = 18.2

Convert pKs to Ks:

pKs = –log(Ks)

Ks = 10–pKs = 10–18.2 = 6.31×10–19


Ag3PO4(s) ⇌ 3Ag+(aq) + PO43–(aq)

___S__________3S_________S_______

Thus,

[Ag+] = 3S

[PO43–] = S


The Ks expression for Ag3PO4(s) is:

Ks = [Ag+]3 × [PO43–]

Therefore,

6.31×10–19 = [3S]3 × [S] = 27S3 × S = 27S4

S4 = (6.31×10–19) / (27) = 2.337×10–20

S = (2.337×10–20)1/4 = 1.236×10–5

[Ag+] = 3S = 3 × (1.236×10–5) = 3.7×10–5

[Ag+] = 3.7×10–5 M


Answer: The concentration of silver (Ag+) in a saturated solution will be 3.7×10–5 M